# Diffraction by a Right-Angle Penetrable Wedge: Closed-Form Solution for $\nu=\sqrt{2}$

### v2

@@ A.1 Scope and provenance of the residue tables
  • the root-of-unity relations $\omega=i$ and $\omega^{4}=1$,
@@ A.2 Orbit reduction and mode matrices
For the right-angle wedge the two faces differ by a rotation of $2\theta_{w}=\pi/2$. On the lemniscatic surface this rotation is implemented by the automorphism $\tau(t,Y)=(it,-Y)$, see §2. After orbit closure one obtains, for each DFT mode $k\in\{0,1,2,3\}$, a matrix Wiener–Hopf/Riemann–Hilbert jump relation of the form
$M_{U,k}(t,Y)\,\widehat{U}_{k}^{+}(t,Y)=M_{V,k}(t,Y)\,\widehat{U}_{k}^{-}(t,Y)+\widehat{H}_{k}(t,Y),\qquad(t,Y)\in\Gamma,$ (86)
where $\Gamma=\{|t|=1\}$ is the physical cut, $\widehat{U}_{k}^{\pm}$ denote the boundary values on the two sides of $\Gamma$, and $\widehat{H}_{k}$ is the DFT forcing term generated by the incident wave. A derivation of such an orbit/DFT reduction for penetrable wedge systems is standard; see, for example, [6].
In the impedance-matched case the face-coupling matrix in Proposition 1.4 depends only on the Snell derivative $w^{\prime}(z)=dw/dz$. In the lemniscatic uniformization one has $w^{\prime}(z)=g^{\prime}(t,Y)$ with
$g^{\prime}(t,Y)=\sqrt{2}\,\frac{t^{2}-1}{Y},\qquad g^{\prime}(\tau(t,Y))=\sqrt{2}\,\frac{t^{2}+1}{Y},$
by (14) and $\tau(t,Y)=(it,-Y)$. Define
$A_{0}$ $:=\frac{1}{2}\bigl(1+g^{\prime}(t,Y)\bigr),$ $B_{0}$ $:=\frac{1}{2}\bigl(1-g^{\prime}(t,Y)\bigr),$
$A_{1}$ $:=\frac{1}{2}\bigl(1-g^{\prime}(\tau(t,Y))\bigr),$ $B_{1}$ $:=\frac{1}{2}\bigl(1+g^{\prime}(\tau(t,Y))\bigr),$
and let $\omega=i$. A convenient normalization of the mode coefficient matrices is
$M_{U,k}=\begin{pmatrix}-\omega^{-k}A_{1}&A_{0}\\[2.0pt] -\omega^{k}B_{1}&B_{0}\end{pmatrix},\qquad M_{V,k}=\begin{pmatrix}-\omega^{-k}B_{1}&B_{0}\\[2.0pt] -\omega^{k}A_{1}&A_{0}\end{pmatrix}$. (87)
$\Delta_{U,k}$ $=\det M_{U,k}=(-\omega^{-k}A_{1})B_{0}-A_{0}(-\omega^{k}B_{1})$ (88)
$=\omega^{k}A_{0}B_{1}-\omega^{-k}A_{1}B_{0},$
$M_{U,k}^{-1}$ $=\frac{1}{\Delta_{U,k}}\begin{pmatrix}B_{0}&-A_{0}\\[2.0pt] \omega^{k}B_{1}&-\omega^{-k}A_{1}\end{pmatrix}$.
$\Delta_{V,k}$ $=\det M_{V,k}=(-\omega^{-k}B_{1})A_{0}-B_{0}(-\omega^{k}A_{1})$ (89)
$=\omega^{k}A_{1}B_{0}-\omega^{-k}A_{0}B_{1},$
$M_{V,k}^{-1}$ $=\frac{1}{\Delta_{V,k}}\begin{pmatrix}A_{0}&-B_{0}\\[2.0pt] \omega^{k}A_{1}&-\omega^{-k}B_{1}\end{pmatrix}$.
@@ A.3 Forcing residues and coefficient extraction
The forcing term $\widehat{H}_{k}$ is meromorphic on $\Sigma_{\mathrm{lem}}$ with simple poles at the forcing set $\{u_{\ell}\}$ defined in §4. Its residues are computed directly from the incident spectral density and the local phase $w_{m}(u)$ along the corresponding orbit branch. In particular, the scalar incident residue
@@
For reproducibility, one may express the forcing residues in the mode variables in the following uniform form. Let $\omega=i$ and define $j=j(\sigma,\varepsilon_{w})$ by
$(+,+)\mapsto 3,\qquad(+,-)\mapsto 1,\qquad(-,+)\mapsto 4,\qquad(-,-)\mapsto 2.$
Let $(A_{m},B_{m})=(A_{0},B_{0})$ for $m$ even and $(A_{m},B_{m})=(A_{1},B_{1})$ for $m$ odd. Then
$\operatorname*{Res}_{u=u_{\ell}}\widehat{H}_{k}(u)=\omega^{-km}r_{I}(\ell)\,v_{k}^{(m,j)},$
where the 22–vector $v_{k}^{(m,j)}$ is given, for $j=1,2,3,4$, by
$\begin{array}[]{ll}j=1:&v_{k}^{(m,1)}=\bigl(\omega^{-k}A_{m},\;\omega^{k}B_{m}\bigr)^{\top},\\[2.84526pt] j=2:&v_{k}^{(m,2)}=\bigl(\omega^{-k}B_{m},\;\omega^{k}A_{m}\bigr)^{\top},\\[2.84526pt] j=3:&v_{k}^{(m,3)}=\bigl(-A_{m},\;-B_{m}\bigr)^{\top},\\[2.84526pt] j=4:&v_{k}^{(m,4)}=\bigl(-B_{m},\;-A_{m}\bigr)^{\top}.\end{array}$
@@
$g_{k}(\ell):=\operatorname*{Res}_{u=u_{\ell}}\widehat{U}_{k}(u)=M_{U,k}(u_{\ell})^{-1}\,\operatorname*{Res}_{u=u_{\ell}}\widehat{H}_{k}(u).$ (90)
@@
where $e_{1}=(1,0)^{\top}$ and $e_{2}=(0,1)^{\top}$.
Substituting (87)–(91) together with the explicit forcing residues above and simplifying using only the lemniscatic relation $Y^{2}=2(t^{4}+1)$ (and $\omega^{4}=1$) yields the closed forms for $\alpha_{\ell}$, $\beta_{\ell}$ and $C_{\ell}$ recorded in Propositions 6.16.3. No additional identities are used.
@@
Fix a forcing label $\ell=(m,\sigma,\varepsilon_{w})$, form $j=j(\sigma,\varepsilon_{w})$, and evaluate (90)–(91) using the forcing residues and the explicit inverses (88). After reducing with $Y^{2}=2(t^{4}+1)$, the resulting expressions for $\alpha_{\ell}$, $\beta_{\ell}$ and $C_{\ell}$ coincide with the corresponding entries in Propositions 6.16.3.
@@
All quantities entering (90)–(91) are rational in $(t,Y)$ and $\omega$ once the lemniscatic curve constraint $Y^{2}=2(t^{4}+1)$ is imposed. For a fixed case $(m\bmod 2,j)$, insert the forcing residue vector $\operatorname*{Res}\widehat{H}_{k}(u_{\ell})$, compute $g_{k}(\ell)=M_{U,k}^{-1}(u_{\ell})\operatorname*{Res}\widehat{H}_{k}(u_{\ell})$, and then read off $t_{\ell}^{4}e_{1}^{\top}g_{1}(\ell)$, $t_{\ell}^{4}e_{2}^{\top}g_{3}(\ell)$ and $\frac{1}{4}\sum_{k=0}^{3}e_{1}^{\top}g_{k}(\ell)$. The subsequent simplification is algebraic and uses only $Y^{2}=2(t^{4}+1)$ and $\omega^{4}=1$. ∎
A.4 Worked examples
We illustrate the residue extraction in two representative cases. Throughout we use only the matrix definitions (87)–(88), the forcing-residue vectors described above, and the lemniscatic identity $Y^{2}=2(t^{4}+1)$.
Example A (even $m$, case $m=0$, $j=1$).
Here $\chi_{0}=1$. For $k=1$ and $m=0$ the forcing residue vector (case $j=1$) is
$\operatorname*{Res}\widehat{H}_{1}(u_{\ell})=r_{I}(\ell)\begin{pmatrix}\omega^{-1}A_{0}\\[1.0pt] \omega B_{0}\end{pmatrix}=r_{I}(\ell)\begin{pmatrix}-iA_{0}\\[1.0pt] iB_{0}\end{pmatrix}$. (92)
Using (88) we obtain
$e_{1}^{\top}g_{1}(\ell)=\frac{1}{\Delta_{U,1}}\bigl(B_{0}(-iA_{0})+(-A_{0})(iB_{0})\bigr)r_{I}(\ell)=\frac{-2iA_{0}B_{0}}{\Delta_{U,1}}\,r_{I}(\ell).$ (93)
Since $\omega=i$ and $\omega^{-1}=-i$, $\Delta_{U,1}=\omega A_{0}B_{1}-\omega^{-1}A_{1}B_{0}=i(A_{0}B_{1}+A_{1}B_{0})$. Substituting the explicit lemniscatic coefficients $A_{0}=\tfrac{1}{2}(1+\sqrt{2}(t^{2}-1)/Y)$, $B_{0}=\tfrac{1}{2}(1-\sqrt{2}(t^{2}-1)/Y)$, $A_{1}=\tfrac{1}{2}(1-\sqrt{2}(t^{2}+1)/Y)$, $B_{1}=\tfrac{1}{2}(1+\sqrt{2}(t^{2}+1)/Y)$ into (93) and reducing with $Y^{2}=2(t^{4}+1)$ yields
$e_{1}^{\top}g_{1}(\ell)=-\frac{r_{I}(\ell)}{t_{\ell}^{2}},\qquad\alpha_{\ell}=t_{\ell}^{4}\,e_{1}^{\top}g_{1}(\ell)=-r_{I}(\ell)t_{\ell}^{2},$ (94)
which is the (meven,j=1)(m\ $\mathrm${even},j=1) entry in Proposition 6.1. A similar computation using (91) gives $C_{\ell}=-r_{I}(\ell)/(2t_{\ell}^{2})$, matching Proposition 6.3.
Example B (odd $m$, case $m=1$, $j=3$).
Here $\psi_{1}=i$ and $\kappa_{1}=-1$. For $k=1$ the forcing residue vector (case $j=3$) is
$\operatorname*{Res}\widehat{H}_{1}(u_{\ell})=\omega^{-1}r_{I}(\ell)\begin{pmatrix}-A_{1}\\[1.0pt] -B_{1}\end{pmatrix}=i\,r_{I}(\ell)\begin{pmatrix}A_{1}\\[1.0pt] B_{1}\end{pmatrix}$. (95)
Applying (88) and simplifying algebraically using $Y^{2}=2(t^{4}+1)$ gives
$e_{1}^{\top}g_{1}(\ell)=-\frac{Y_{\ell}}{\sqrt{2}\,t_{\ell}^{2}}\,r_{I}(\ell),\qquad\alpha_{\ell}=t_{\ell}^{4}\,e_{1}^{\top}g_{1}(\ell)=\kappa_{1}\,r_{I}(\ell)\,\frac{Y_{\ell}}{\sqrt{2}}\,t_{\ell}^{2},$ (96)
which coincides with the (modd,j=3)(m\ $\mathrm${odd},j=3) entry in Proposition 6.1. The remaining cases, including the $\beta_{\ell}$ and $C_{\ell}$ tables, follow from the same residue-insertion step combined with algebraic reduction on $\Sigma_{\mathrm{lem}}$.
A.5 Remark on symmetries
The case structure in Propositions 6.16.3 reflects the symmetries of the lemniscatic surface. The automorphism $\tau$ permutes orbit indices $m\mapsto m+1$ and exchanges the roles of the polynomials $(t^{2}-1)$ and $(t^{2}+1)$, while the involution $\eta(t,Y)=(t,-Y)$ flips the branch of $Y$. These symmetries reduce the residue computation to a small number of representative configurations.

### v3

@@ A.1 Scope and provenance of the residue tables
  • the root-of-unity relations $\omega=\mathrm{i}$ and $\omega^{4}=1$,
@@ A.2 Orbit reduction and mode matrices
For the right-angle wedge the two faces differ by a rotation of $2\theta_{w}=\pi/2$. On the lemniscatic surface this rotation is implemented by the automorphism $\tau(t,Y)=(\mathrm{i}t,-Y)$, see §2. After orbit closure one obtains, for each DFT mode $k\in\{0,1,2,3\}$, a matrix Wiener–Hopf/Riemann–Hilbert jump relation of the form
$M_{U,k}(t,Y)\,U_{k}^{b,+}(t,Y)=M_{V,k}(t,Y)\,U_{k}^{b,-}(t,Y)+H_{k}^{b}(t,Y),\qquad(t,Y)\in\Gamma,$ (72)
where $\Gamma=\{|t|=1\}$ is the physical cut, $U_{k}^{b,\pm}$ denote the boundary values on the two sides of $\Gamma$, and $H_{k}^{b}$ is the DFT forcing term generated by the incident wave. A derivation of such an orbit/DFT reduction for penetrable wedge systems is standard; see, for example, [6].
In the impedance-matched case the face-coupling matrix in Proposition 1.4 depends only on the Snell derivative $w^{\prime}(z)=\,\mathrm{d}w/\,\mathrm{d}z$. In the lemniscatic formulation one has $w^{\prime}(z)=g^{\prime}(t,Y)$ with
$g^{\prime}(t,Y)=\frac{\sqrt{2}(t^{2}-1)}{Y},\qquad g^{\prime}(\tau(t,Y))=\frac{\sqrt{2}(t^{2}+1)}{Y},$
by (17) and $\tau(t,Y)=(\mathrm{i}t,-Y)$. Define
$A_{0}$ $=\frac{1}{2}\bigl(1+g^{\prime}(t,Y)\bigr),$ $B_{0}$ $=\frac{1}{2}\bigl(1-g^{\prime}(t,Y)\bigr),$
$A_{1}$ $=\frac{1}{2}\bigl(1-g^{\prime}(\tau(t,Y))\bigr),$ $B_{1}$ $=\frac{1}{2}\bigl(1+g^{\prime}(\tau(t,Y))\bigr).$
and let $\omega=\mathrm{i}$. A convenient normalization of the mode coefficient matrices is
$M_{U,k}=\begin{pmatrix}-\omega^{-k}A_{1}&A_{0}\\ -\omega^{k}B_{1}&B_{0}\end{pmatrix},\qquad M_{V,k}=\begin{pmatrix}-\omega^{-k}B_{1}&B_{0}\\ -\omega^{k}A_{1}&A_{0}\end{pmatrix}$. (73)
Write $\Delta_{U,k}:=\det M_{U,k}=\omega^{k}A_{0}B_{1}-\omega^{-k}A_{1}B_{0}$. Then
$M_{U,k}^{-1}=\frac{1}{\Delta_{U,k}}\begin{pmatrix}B_{0}&-A_{0}\\ \omega^{k}B_{1}&-\omega^{-k}A_{1}\end{pmatrix}$. (74)
Similarly $\Delta_{V,k}=\omega^{k}A_{1}B_{0}-\omega^{-k}A_{0}B_{1}$ and
$M_{V,k}^{-1}=\frac{1}{\Delta_{V,k}}\begin{pmatrix}A_{0}&-B_{0}\\ \omega^{k}A_{1}&-\omega^{-k}B_{1}\end{pmatrix}$. (75)
@@ A.3 Forcing residues and coefficient extraction
The forcing term $H_{k}^{b}$ is meromorphic on $\Sigma_{\mathrm{lem}}$ with simple poles at the forcing set $\{u_{\ell}\}$ defined in §4. Its residues are computed directly from the incident spectral density and the local phase $w_{m}(u)$ along the corresponding orbit branch. In particular, the scalar incident residue
@@
For reproducibility, one may express the forcing residues in the mode variables in the following uniform form. Let $\omega=\mathrm{i}$ and define $j=j(\sigma,\varepsilon_{w})$ by (24). Let $(A_{m},B_{m})=(A_{0},B_{0})$ for $m$ even and $(A_{m},B_{m})=(A_{1},B_{1})$ for $m$ odd. Then
$\operatorname*{Res}_{u=u_{\ell}}H_{k}^{b}(u)=\omega^{-km}r_{I}(\ell)\,v_{k}^{(m,j)},$
where the $2$-vector $v_{k}^{(m,j)}$ is given, for $j=1,2,3,4$, by
$j=1:$ $\quad v_{k}^{(m,1)}=\binom{\omega^{-k}A_{m}}{\omega^{k}B_{m}}$, $j=2:$ $\quad v_{k}^{(m,2)}=\binom{\omega^{-k}B_{m}}{\omega^{k}A_{m}}$,
$j=3:$ $\quad v_{k}^{(m,3)}=\binom{-A_{m}}{-B_{m}}$, $j=4:$ $\quad v_{k}^{(m,4)}=\binom{-B_{m}}{-A_{m}}$.
@@
$g_{k}(\ell):=\operatorname*{Res}_{u=u_{\ell}}U_{k}^{b}(u)=M_{U,k}(u_{\ell})^{-1}\,\operatorname*{Res}_{u=u_{\ell}}H_{k}^{b}(u).$ (76)
@@
where $e_{1}=(1,0)^{\top}$ and $e_{2}=(0,1)^{\top}$. Substituting (73)–(77) together with the explicit forcing residues above and simplifying using only the lemniscatic relation $Y^{2}=2(t^{4}+1)$ (and $\omega^{4}=1$) yields the closed forms recorded in Propositions 6.16.3.
@@
Fix a forcing label $\ell=(m,\sigma,\varepsilon_{w})$, form $j=j(\sigma,\varepsilon_{w})$, and evaluate (76)–(77) using the forcing residues and the explicit inverse (74). After reducing with $Y^{2}=2(t^{4}+1)$, the resulting expressions for $\alpha_{\ell},\beta_{\ell},C_{\ell}$ coincide with the corresponding entries in Propositions 6.16.3.
@@
All quantities entering (76)–(77) are rational in $(t,Y)$ and $\omega$ once the lemniscatic curve constraint $Y^{2}=2(t^{4}+1)$ is imposed. For a fixed case ($m\bmod 2$, $j$), insert the forcing residue vector $\operatorname*{Res}H_{k}^{b}(u_{\ell})$, compute $g_{k}(\ell)=M_{U,k}^{-1}(u_{\ell})\operatorname*{Res}H_{k}^{b}(u_{\ell})$, and then read off the linear functionals in (77). The subsequent simplification is algebraic and uses only $Y^{2}=2(t^{4}+1)$ and $\omega^{4}=1$. ∎